OpenStudy (anonymous):

ln(−10 − x) + ln 3 = ln(2 − x)

OpenStudy (freckles):

use ln(a)+ln(b)=ln(ab) on the left hand side

OpenStudy (anonymous):

ok i have it

OpenStudy (anonymous):

0,2

OpenStudy (anonymous):

the answer is x= 0,2?

OpenStudy (anonymous):

\[\ln (-10-x)+\ln 3=\ln (2-x)\]\[\ln [(-10-x)(3)]=\ln (2-x)\]\[\ln (-30-3x)=\ln (2-x)\]\[\log_{e} \ln (-30-3x)=\log_{e} \ln (2-x)\]\[-30-3x=2-x\]\[-32=2x\]\[-16=x\]

OpenStudy (anonymous):

no sorry the drawing is the answer